OpenStudy (anonymous):

log(4x+1)=1+log(x-1)

OpenStudy (anonymous):

Combine like terms with each others : Log(4x+1)-Log(x-1)=1 Then Use Log property Can you continue ?

OpenStudy (anonymous):

the left hand side becomes \[\log\left(\frac{4x+1}{x-1}\right)=1\]

OpenStudy (anonymous):

Then change it to exponential

OpenStudy (anonymous):

if you cant put a negative in the log what happens to the left hand side? or does x-1 become x+1

OpenStudy (anonymous):

After you change it to exponential ,it will be 10^1=(4x+1)/(x-1) multiply by (x-1) to both sides and solve for x

OpenStudy (anonymous):

There is a variable in log (x-1) in order to avoid being a negative though

OpenStudy (anonymous):

10x-10 = 4x+1 10x = 4x + 11 6x = 11 x = 11/6 i feel like thats not right lol

OpenStudy (anonymous):

\[\Huge \color{Red}{That's~Right~♥ ❤ ❥}\]

OpenStudy (anonymous):

OMG :D yay :) so is it most proper to write as a fraction or as a decimal

OpenStudy (anonymous):

it says "type an exact answer using radicals or log functions as needed"

OpenStudy (anonymous):

I think fraction would be better :P